Water stands at a depth of 15 m behind a reservoir dam. A horizontal pipe 4 cm in diameter passes through the dam 6 m below the surface of water as shown. There is plug which secures the pipe opening. Then the friction between the plug and pipe wall is:

Grade:12

3 Answers

Chetan Mandayam Nayakar
13 years ago

pressure=dgh, d=1000kg/m^3, h=6m, g=9.8m/s^2, area=pi*r2,r=2cm,friction force=pressure*area

Neha
3 years ago
Friction=force applied 
So, force applied= area×pressure 
R= d\2= 0.04/2=0.02m
Area= pie× R*2=3.14×0.02×0.02=0.001256m*2
And pressure=P2- P1=P0+ rho×g×h - P0= rho× g× h= 1000×10×6=60000
Therefore, force applied= friction = area× pressure
                                           =0.001256×60000
                                            =75.36N
 
Vikas TU
3 years ago
Dear student 
Above ans are incorrect 
P = rho * g*h 
= 58800N/m^2
Splug = pi*R^2 
Magnitude of the frictional force b/n plug and pipe wall
F = S* P = 295.41N 
V = sqrt(2*g*h)
=7.67 m/s